# When is Multicalibration Post-Processing Necessary?

### v1

@@ 5.1 Limitations and Conclusion
Another limitation of our results is that they are restricted to binary classification problems. While multicalibration algorithms do extend to multiclass problems, this extension comes at a severe cost of sample efficiency usually exponential in the number of labels (Zhao et al., , 2021). We show that — at least for tabular datasets — current multicalibration algorithms do not significantly improve upon a competitive ERM. If we were to further burden the multicalibration algorithm with the larger sample complexity of an additional label, we do not expect that their performance will improve. Nonetheless, we plan to investigate the multiclass setting in future work, and believe that those findings will be consistent with the results present in this paper.
@@ Acknowledgements.
SD was supported by the Department of Defense through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. This work was also supported by NSF CAREER Award CCF-2239265 and an Amazon Research Award. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of sponsors such as Amazon or NSF. The authors would like to thank Bhavya Vasudeva for discussions that were helpful in the design of early experiments, and Eric Zhao for help in utilizing the HJZ algorithms.
@@ Appendix A Additional Related Work
Subgroup Robustness. Several works in subgroup robustness literature examine the performance of NNs by worst-group-accuracy, particularly in cases where NNs tend to rely on spurious correlations. Recent works by Kirichenko et al., (2023); LaBonte et al., (2023) propose last-layer fine-tuning as a simple and computationally inexpensive way to do exactly that. Indeed, Mao et al., (2023) extend the method to address fairness concerns, appending a fairness constraint to the training objective during fine-tuning. Such works examine only one sensitive attribute at a time, and often consider the disjoint groups produced by unique values of this attribute in conjunction with label.
A.1 Equivalence of Early Stopping and Hyperparameter Sweeps
@@ C.4 Groups for Tabular Datasets
C.5 Groups for Image Datasets
@@
C.6 Groups for Language Datasets
@@
C.7 Dataset Usage and Licensing

### v2

@@ 7 Experimental Limitations and Conclusion
Another limitation of our results is that they are restricted to binary classification problems. While multicalibration algorithms do extend to multiclass problems, this extension comes at a severe cost of sample efficiency usually exponential in the number of labels (Zhao et al., , 2021). We show that — at least for tabular datasets — current multicalibration algorithms do not significantly improve upon a competitive and calibrated ERM baseline. If we were to further burden the multicalibration algorithm with the larger sample complexity of an additional label, we do not expect that their performance will improve. Nonetheless, we plan to investigate the multiclass setting in future work, and believe that those findings will be consistent with the results present in this paper.
Finally, we do not offer much explanation of why we see differing performance of the two algorithms HJZ and HKRR; we offer some discussion of this in Section A.3, but more is warranted in future work.
@@ Acknowledgements.
SD was supported by the Department of Defense through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. This work was also supported by NSF CAREER Award CCF-2239265 and an Amazon Research Award. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of sponsors such as Amazon or NSF. The authors would like to thank Bhavya Vasudeva for discussions that were helpful in the design of early experiments, and Eric Zhao for help in utilizing the HJZ algorithms. The authors also sincerely thank the anonymous Neurips reviewers for providing detailed feedback and discussion which greatly improved and clarified parts of this work.
@@ Appendix A Additional Related Work
Subgroup Robustness. Several works in subgroup robustness literature examine the performance of NNs by worst-group-accuracy, particularly in cases where NNs tend to rely on spurious correlations. Recent works by Kirichenko et al., (2023); LaBonte et al., (2023) propose last-layer fine-tuning as a simple and computationally inexpensive way to do exactly that. Indeed, Mao et al., (2023) extend the method to address fairness concerns, appending a fairness constraint to the training objective during fine-tuning. Such works examine only one sensitive attribute at a time, and often consider the disjoint groups produced by unique values of this attribute in conjunction with label. Rosenfeld and Garg, (2023) show connections between robustness and distribution shift for neural networks via unlabeled test data.
In general, we view multi-group robustness as a notion of robustness which, like multicalibration, aims to “respect” multiple groups simultaneously. Perhaps the closest connection between multigroup robustness and multicalibration is: they intuitively may share similar mechanisms (as the reviewer noted). That is, the mechanistic question in both cases is to understand which groups can be easily computed from the predictor’s "features" – ie, which groups are easy for the predictor to distinguish.
A.1 Further Discussion of Błasiok et al., (2024)
The connection to our work is subtle. In Błasiok et al., (2024), the authors consider performing ERM over some $\mathcal{C}^{\prime}\supseteq\mathcal{C}$ so that the resulting model is multicalibrated with respect to $\mathcal{C}$. In their case, $\mathcal{C}^{\prime}$ is a family of neural networks (NNs), and $\mathcal{C}$ is some family of smaller NNs. In many of our experiments, we indeed perform (approximate) ERM over some family of NNs, and one might expect this to result in multicalibration with respect to our finite collection of groups $G$, which are easily computable by some class of smaller NNs. This is because for NNs, it is possible to represent arbitrarily-complex groups by simply taking "large enough" networks, without making design choices specific to the group structure.
In our tabular experiments, we restrict our experiments to small multi-layer perceptron networks of 3-4 layers, whose smaller subnetworks may not be learning complex functions capturing groups of interest (see Appendix E for model descriptions). Nonetheless, we find that these models possess latent multicalibration properties, as discussed further in Observation 1 in Section 3. Our vision and language experiments show that post-processing does have positive impact, suggesting that the models sub-networks are not sufficiently capturing the groups of interest (see Section 4).
A.2 Equivalence of Early Stopping and Hyperparameter Sweeps
@@
A.3 Discussion on Dynamics and Performance of HJZ and HKRR
The dynamics of both of the algorithms HJZ and HKRR are complex. In particular, the performance of the algorithms depend on at least the following parameters:
  1. 1.
    Distribution of initial predictions output by the models (i.e. input to post-processing algorithm);
  2. 2.
    Choice of hyperparameters for HKRR and HJZ;
  3. 3.
    "Complexity” or “expressiveness” of the groups; and
  4. 4.
    Number of available samples, and whether the samples are re-used from training or not.
In our work, we focus mainly on (2) and (4). We discuss (3) to an extent in Section 6.1, but teasing apart exactly how (1) and (3) contribute to the performance of multicalibration post-processign algorithms is certainly an interesting avenue for future work. We believe that part of the reason for the superiority of HKRR in language/vision data may be explained within the lens of (2) and (4). That is, we may have found better hyperparameters for HKRR with a wider search, and the sample complexity may be better in practice than the game-theoretic approach offered by HJZ.
Due to computational constraints and the added dimension of choosing how much data to save for calibration, we search a large — but not all-encompassing — collection of hyperparameters for each of the multicalibration algorithms tested. With regards to dataset size, 3 of the 4 vision/language datasets are noticeably larger than the tabular datasets (by at least 100k samples). It is possible that HKRR generally performs better on such dataset sizes, or that the optimal hyperparameters for HJZ change significantly in this larger-sample regime.
Understanding why HJZ may be more stable to hyperparameter choices (see Observation 5 in Section 3) is a more challenging question to answer, since it likely has to do with internal game dynamics in the learning algorithm. In particular, by choosing various online learning algorithms, HJZ implements a family of multicalibration post-processing methods. We test all algorithms from this family with varying parameters. It is possible that the family of algorithms itself somehow has a shift in stability as we scale to a large data regime (4). However, as analyzing even a singular algorithm (e.g. HKRR) is challenging, we are not sure that speculating about the stability of a family of algorithms is currently possible, and hence, leave this to future work.
@@ C.4 Groups for Tabular Datasets
C.5 Alternate Groups for Tabular Datasets
To validate our observations on tabular data, we repeated all experiments on each tabular dataset with an alternate collection of groups. Like the original groups, these groups are defined by a feature or conjunction of two features. We provide the alternate group definitions here, and present results on these groups in Appendix G.
{adjustwidth}
-1in-1in group name n samples fraction y mean Associates Degree Male 7331 0.0375 0.4957 Associates Degree Female 8372 0.0428 0.3186 Divorced Female 10652 0.0544 0.4415 Under Part Time 16525 0.0845 0.1025 Part Time 55269 0.2825 0.1408 Full Time 135989 0.6950 0.5214 Over Full Time 11471 0.0586 0.6441 Not White 74659 0.3816 0.3574 Government Employee 29121 0.1488 0.5337 Private Employee 166544 0.8512 0.3890 Under 21 10166 0.0520 0.0106 Middle Aged 81582 0.4169 0.5064 Dataset 195665 1.0000 0.4106
Figure 11: ACS Income alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Job $=$ Management, Age $<$ 50 7091 0.1568 0.1393 Job $=$ Technician, Age $<$ 30 436 0.0096 0.1858 Job $=$ Blue-Collar, Age $>$ 50 1075 0.0238 0.0679 Married, Education $=$ Primary 5246 0.1160 0.0755 Single, Education $=$ Tertiary 4792 0.1060 0.1836 Housing $=$ Yes, Age $<$ 30 1621 0.0359 0.0993 Housing $=$ No, Age $<$ 30 1429 0.0316 0.3037 Under 21 305 0.0067 0.3115 Middle Age 23841 0.5273 0.0977 Senior Age 961 0.0213 0.4225 Dataset 45211 1.0000 0.1170
Figure 12: Bank Marketing alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Single, Male 6553 0.2184 0.2266 Single, Female 15964 0.5321 0.2093 Young Adult 9618 0.3206 0.2284 Middle Aged 8872 0.2957 0.2356 Education $=$ High School, Female 2927 0.0976 0.2364 Education $=$ University, Female 8656 0.2885 0.2220 Education $=$ High School, Single 1909 0.0636 0.2368 Education $=$ High School, Married 2861 0.0954 0.2635 Education $=$ University, Single 7020 0.2340 0.2306 Education $=$ University, Married 6842 0.2281 0.2435 Education $=$ Graduate, Single 6809 0.2270 0.1842 Dataset 30000 1.0000 0.2212
Figure 13: Credit Default alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Loan Type 1 87857 0.7694 0.7327 Loan Type 2 17587 0.1540 0.8125 Loan Type 3 8047 0.0705 0.8311 HUD Median Family Income $>$ 50k 107450 0.9410 0.7583 HUD Median Family Income $\leq$ 50k 6735 0.0590 0.6578 Has Co-Applicant 42535 0.3725 0.8079 Agency $=$ OCC 5966 0.0522 0.8235 Agency $=$ FRS 3879 0.0340 0.8729 Agency $=$ FDIC 6951 0.0609 0.8708 Agency $=$ NCUA 10626 0.0931 0.6882 Agency $=$ HUD 64915 0.5685 0.7562 Agency $=$ CFPB 21848 0.1913 0.6939 Loan Type $=$ 1 to 4 Family 87857 0.7694 0.7327 Loan Type $=$ Manufactured Housing 17587 0.1540 0.8125 Loan Type $=$ Multi-Family 8047 0.0705 0.8311 Dataset 114185 1.0000 0.7524
Figure 14: HMDA alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Under 21 3772 0.3405 0.0607 Middle Age 2874 0.2594 0.1990 Senior Age 1304 0.1177 0.4862 Sex $=$ 1 5281 0.4767 0.1274 Sex $=$ 2 5798 0.5233 0.2077 White 3958 0.3573 0.2534 Active Duty Group 2 6454 0.5825 0.1432 Marriage Group 1 3645 0.3290 0.2222 Marriage Group 2 450 0.0406 0.4867 Pregnancy Group 1 124 0.0112 0.4274 Pregnancy Group 2 2167 0.1956 0.1398 Insurance Group 1 5926 0.5349 0.1790 Insurance Group 2 3890 0.3511 0.1979 Dataset 11079 1.0000 0.1694
Figure 15: MEPS alternate groups.
C.6 Groups for Image Datasets
@@
C.7 Groups for Language Datasets
@@
C.8 Dataset Usage and Licensing